Can someone show me step by step the correct answer. Thanyou It will be greatly appreciated.
Question 1
The following represents the mean scores on the same set of 5 statistics exams given in Spring 2004 and Fall 2004. Test the claim that the mean scores for the Spring 2004 semester were greater than or equal to the mean scores for the Fall 2004 semester. Use a= 0.10
Spring 2004
Test 1 =46.0, Test 2= 47.1, Test 3= 55.7, Test 4= 58.2, Test 5= 62.4

Fall 2004
Test 1= 46.3, Test 2= 47.2, Test 3= 55.2, Test 4= 57.6, Test 5=63.6


Answer by stanbon(75887)

The following represents the mean scores on the same set of 5 statistics exams given in Spring 2004 and Fall 2004. Test the claim that the mean scores for the Spring 2004 semester were greater than or equal to the mean scores for the Fall 2004 semester. Use a= 0.10
Spring 2004
Test 1 =46.0, Test 2= 47.1, Test 3= 55.7, Test 4= 58.2, Test 5= 62.4
Fall 2004
Test 1= 46.3, Test 2= 47.2, Test 3= 55.2, Test 4= 57.6, Test 5=63.6

-----------------
Find the mean of the Spring scores and the mean of the Fall scores:
Spring: 53.88; Fall:53.98
---------
Ho: mu1-mu2=0
Ha: ma1-mu2>0
t(53.98-53.88) = 0.11/sqrt((7.12)^2/5 + (7.23)^2/5)= 0.11/4.54=0.024
Critical value for alpha = 0.10 = ?
Comment: I don't have a t-table available to find the critical value.
But using a TI calculator I get a P-value 0f 0.5085 so we can
Fail to reject Ho.
Conclusion: The means are not different.
==============
Cheers,
Stan H.